Shift report 20010916 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Sep 16 2001 - 16:10:36 EDT

  • Next message: Flemming Videbaek: "shift list"

    ------------------------------------------------------------
    Date: Sun Sep 16 16:10:35 2001, Shift: 08:00-16:00
    Supervisor: JH Lee, with: 
    Log book#8 pages 139 - 140
    
    Shift Summary:
    ------------------------------------------------------------
     Beam:
      - Cogged beam from the previous shift (since 3:30) 
      - At 9:45 vote for dump (~at ZDC: 20-25).
        PHENIX voted for dump earlier, PHOBOS and BRAHMS
        voted at the almost same time
      - Beam dump by Blue Quench Interlock at 10:10 
      - No beam since then  
    
     Data: 
      - MRS at 90 (2A) FS at 8 (1/5 B)
      - Run 4757 with 164K trigger1 events
      
     Detector:
      - No New Detector Problems
    
     Comments: 
      - Please make sure that the Fast Electronics Hut door 
        is closed and locked by pushing it hard. It might 
        need to be fixed.      
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Sep 16 2001 - 16:11:06 EDT